# Phase-field modeling of radiation-induced composition redistribution: An application to additively manufactured austenitic Fe-Cr-Ni

### v1

@@ 3 Parameterization
$P_{\nu}$ Point defect production rate $2\times 10^{-6}$ dpa/s
@@
$S^{f}_{V}$ Vacancy formation entropy 2 $k_{B}$ [11]
$Q_{k}$ Activation energy for vacancy diffusion of $k$ Function [27]
@@ 3.1 Transport properties
Following Yang et al. [27], the Onsager coefficients are calculated from the partial diffusivities $d_{k\nu}$ by using Manning’s relations for concentrated multicomponent alloys [56]:
$L^{\nu}_{kl}=\left(\frac{c_{k}c_{\nu}d_{k\nu}}{RT/V_{m}}\right)\left(\delta_{kl}+\frac{2c_{l}d_{l\nu}}{M_{o}\sum_{j}c_{j}d_{j\nu}}\right),$ (18)
@@
Since the vacancy activation energies $Q_{k}$ are derived from experimental tracer diffusivity data, the vacancy migration energy for diffusion is given by $E^{m}_{Vk}=Q_{k}-E^{f}_{V}$. $Q_{k}$ is found to lie in the 2.84–2.97 eV range, and $E^{f}_{V}=1.8$ eV is chosen based on a first-principles study on dilute FCC Fe-Cr-Ni [58]. This yields an effective migration free energy in the range of 1.2 to 1.25 eV at 500 C. With this approach, $E^{f}_{V}$ is an uncertain parameter that must be more accurately estimated for a concentrated austenitic SS composition. The vacancy diffusivity at 500 C is $D_{V}\approx 5.5\times 10^{5}$ nm2/s and the partial diffusion coefficient ratios are $d_{CrV}/d_{FeV}=1.62$ and $d_{NiV}/d_{FeV}=0.67$. These ratios reflect the faster diffusivity of Cr and the slower diffusivity of Ni in FCC Fe. The SIA parameters in the model are quite uncertain, since they are not experimentally accessible and are challenging to determine via atomistic techniques. $E^{m}_{I}$ = 0.9 eV, as chosen by Yang et al. [27], resulted in a low diffusivity for SIA in comparison to the vacancy. Therefore, we chose a lower value of $E^{m}_{I}=0.6$ eV [27], yielding $D_{I}\approx 1.3\times 10^{7}$ nm2/s at 500 C—an order of magnitude higher than $D_{V}$. With regards to SIA binding, first-principles studies of SIA in dilute [58] and concentrated [59] FCC Fe–Ni–Cr indicate an attractive interaction between SIAs and Cr, a repulsive interaction with Ni, and an ideal interaction with Fe. Thus, the positive value for $E^{b}_{CrI}-E^{b}_{FeI}$ models favorable transport of Cr via SIAs, whereas the negative value for $E^{b}_{NiI}-E^{b}_{FeI}$ models unfavorable transport of Ni via SIAs. The resulting binding factors for the SIA partial diffusivities are ${\beta}_{CrI}=1.37$, ${\beta}_{NiI}=0.84$, and ${\beta}_{FeI}=0.94$. The ratios of SIA partial diffusion coefficients are $d_{CrI}/d_{FeI}=1.63$ and $d_{NiI}/d_{FeI}=0.89$. At lower temperatures ($<500^{\circ}$ C), the increased importance of Cr-SIA binding means that Cr transport via SIA is expected to become more important (i.e. $d_{CrI}/d_{FeI}>d_{CrV}/d_{FeV}$). All simulation of RIS performed in this paper employ the preferential solute-SIA binding discussed above. A comparison of these results with the case without solute-SIA binding is performed in Sec. S6 of the Supplementary Material.
@@ 3.3 Point defect reaction rates
The reaction rate coefficients in Eq. 12 and 14 are defined as follows. We set $P_{V}=P_{I}=2\times 10^{-6}$ dpa/s representing conditions of neutron or proton irradiation [23]. For the present work, we do not consider the formation and effects of clusters in the collision cascade and thus neglect the resulting point defect production bias. The recombination reaction coefficient is given by $R_{VI}=4\pi r_{\circ}(D_{V}+D_{I})/V_{a}$, with $r_{\circ}$ being the recombination radius. The point defect diffusivities are derived from the partial diffusivities in Sec. 3.1 as $D_{\nu}=(c_{Ni}d_{Ni\nu}+c_{Cr}d_{Cr\nu}+c_{Fe}d_{Fe\nu})/f_{\circ}$.
The sink strength of dislocation is given by $k^{2}_{\nu,b}=\rho_{b}Z_{\nu}$, where $Z_{\nu}$ is the sink efficiency of dislocations for absorption of point defects. Realistically, $Z_{\nu}$ is a function of the dislocation density and character [63]. Relations for $Z_{\nu}$ based on calculations of the elastic interaction between dislocation configurations and the point defect have revealed a bias for SIA absorption (i.e., $Z_{I}>Z_{V}$), and this bias increases with dislocation density. For the sake of simplicity, we choose dislocation-density-independent values of $Z_{I}=1.2$ and $Z_{V}=1$, amounting to a bias factor $(Z_{I}-Z_{V})/Z_{V}$ of 20% [54]. We note that estimates of bias factors ranging from 2 to 30% have been reported using different modeling techniques and the discrepancy between these have been discussed in Ref. [64]. The choice of $Z_{I}=1.2$ therefore represents a strong bias.
@@ 3.4 Equilibrium PF properties
For the AM microstructure, per the literature [9, 10], a CW width of $\delta=100$ nm is considered. We choose a dislocation density of $\rho_{b,c}=10^{14}$ m-2 in the cell, and set $\rho_{b,w}$ in the CW to be 5–15 times higher. For $\rho_{b,w}=10^{15}$ m-2 distributed uniformly across the CW width, the dislocation energy of the CW can be estimated as $\frac{1}{2}Gb^{2}\rho_{c,w}\delta\approx 1$ J/m2, where $G=77$ GPa is the shear modulus. To realize a stationary CW of $\delta=100$ nm, we set the PF model parameters as $m_{0}=7.5\times 10^{6}$ J/m3 and $\kappa=7.5\times 10^{-8}$ J/m from Sec. S5 of the Supplementary Material. The interpolation scheme to determine the total atomic thermodynamic factors in Eq. 12 and 14 from the phase-specific atomic thermodynamic factors obtained above is provided in Sec. S2 of the Supplementary Material. The thermodynamic factors for the point defect species are simply given by ${\theta}_{\nu\nu}=(RT/V_{m})/c_{\nu}$.
@@ 4 Implementation
The model is implemented using the open-source MOOSE (Multiphysics Object Oriented Simulation Environment) framework. Weak forms of the partial differential equations are spatially discretized via the finite element method. For time integration, the implicit second-order backward differentiation method is used to obtain the system of nonlinear equations at each time step. These are then solved using Newton’s method. Mesh elements with linear Lagrange shape functions are used for the nonlinear and auxiliary variables. 2D domains are meshed with four-node quadrilateral elements. For both 1D and 2D simulations, a uniform mesh with adaptive time stepping, a nonlinear relative tolerance of $10^{-8}$, and a nonlinear absolute tolerance of $10^{-10}$ are employed. The order parameters for the 2D polycrystal and 2D CW microstructures (Fig. 2) are generated using the Voronoi tessellation and grain growth algorithm implemented in MOOSE [66].
@@ 4.1 GB segregation
Simulations of GB segregation are performed in 1D and 2D with periodic boundary conditions. For the PF implementations, Eqs. 1212, and 13 are solved. For a GB width of $\delta=1$ nm, a mesh element size of $0.1$ nm is chosen, based on a mesh convergence test. The different PF simulation methods adopted in this work are as follows:
@@
In addition to the above PF simulations, we performed sharp interface simulations (see A and B) for RIS at GBs by employing Eq. 12 with boundary conditions. For this, we simply omit the PF sink term $-k^{2}_{g}D_{\nu}(c_{\nu}-c^{e}_{\nu})g_{\text{sink}}$ from Eq. 12 and impose a Dirichlet boundary condition (DBC) or Robin boundary condition (RBC) for the point defect concentration variables at the edges of the domain defining the GB. The DBC sets $c_{\nu}=c^{e}_{\nu}$ at the GB, thus describing an ideal sink. The RBC is applied as a flux normal to the GB given by $J_{\nu}.\hat{n}=D_{\nu}\alpha^{-1}(c_{\nu}-c^{e}_{\nu})$, where $\alpha$ is a characteristic length for point defect absorption  [65]. The effect of this parameter on the non-ideal GB sink behavior is studied in  B. The different sharp-interface simulations of RIS at GBs performed in this work are:
@@ 5 Results
We first present the results of GB segregation in 1D and 2D systems as a function of irradiation dose and temperature. We examine four different GB segregation mechanisms: RIS without TS, TS without RIS, radiation-enhanced TS (RETS) without RIS, and RIS in combination with RETS (RIS+RETS). The effects of grain size and dislocation bias on RIS are then presented. The results obtained through verification and comparison of the 1D PF RIS simulations against sharp-interface simulations are provided in A and B. Next, we present the results of RIS at dislocation CWs in 1D and 2D cells. The effects of cell size, CW density, initial microsegregation, and GB sink are examined.
@@ 5.1.1 Segregation profiles
Concentration profiles for the RIS mechanism are plotted for different doses in Fig. 3(a). Ni enrichment (inverted “V”-shaped profile) and Cr depletion (“V”-shaped profile) are observed, with wide segregation widths of over 100 nm on either side of the GB center. Significant changes to RIS were found to occur between 0.01 dpa (1 hour) and 10 dpa (60 days). Concentration profiles for the RETS mechanism are plotted in Fig. 3(b). Unlike RIS, RETS is seen to evolve rapidly and reach equilibrium at just after 0.01 dpa (1 hour). Moreover, Ni depletion and Cr enrichment, which are opposite to their behaviors during RIS, are observed. In further contrast, RETS shows sharp profiles that only extend to about 0.5 nm on either side of the GB center. Next, Fig. 4 shows profiles from the combined RIS+RETS simulation. Here, two different cases are shown, with TS being strong ($\sigma=0.8$) in one case and moderate ($\sigma=0.9$) in the other. While the low-magnification plots of the profiles (not shown) resemble the monotonic RIS profiles, the high-magnification plots (Figs. 4(a) and 4(b)) clearly show non-monotonic variation close to the GB. For the case with strong TS, a slight Cr enrichment persists at the GB center, even at 10 dpa. The profiles resemble a “W” shape for Cr and an “M” shape for Ni. The non-monotonic profiles are less pronounced for the case of weak TS. Note that for the sake of simplicity, the local GB sink strength $k^{2}_{g}$ was kept the same and only $\sigma$ was varied; however, in a more realistic scenario, both $k^{2}_{g}$ and $\sigma$ would vary as a function of the GB structure.
@@ 5.1.2 Dose and temperature dependence
However, TS without irradiation shows lower segregation at the end of the simulated 60 days than that observed with RETS below 450 C, indicating that equilibrium TS is not achieved during that time frame due to the extremely slow diffusivity at lower temperatures. In contrast, RIS varies strongly with temperature and peaks at around 500 C. Negligible RIS is observed above 750 C due to thermal back diffusion. Conversely, below 500 C, increasing contributions from recombination and preferential SIA diffusivity (see Sec. S6 in the Supplementary Material) lead to decreasing RIS (here, SIAs are mobile but vacancies have reduced mobility; thus, recombination is promoted, resulting in a reduced loss of vacancies to sinks). Compared to RIS alone, RIS+RETS for $\sigma=0.8$ (strong TS) shows lower Ni enrichment at intermediate temperatures, while a transition to Ni depletion is observed below 400 C and above 650 C. For RIS+RETS of Cr, RETS dominates across the temperature range, resulting in an effective Cr enrichment at the GB center. However, at 0.5 nm from the GB center, there is negligible influence from RETS; therefore, the temperature trend follows that of RIS. Finally, we note that for $\sigma=0.95$ (weak TS, not shown), the RIS+RETS results are very close to that of RIS, even at the GB center.
@@ 6.1 PF model for GB segregation
A multi-order-parameter PF model was developed that describes both RIS and RETS in multicomponent polycrystals. The formulation also preserves the complete set of Onsager coefficients from the Onsager force-flux relations, thus allowing non-ideal kinetics ($L_{kV}/L^{\nu}_{kk}\neq-1$) to be captured. Therefore, RIS mechanisms of solute-vacancy exchange or inverse Kirkendall effect ($L_{kV}/L^{V}_{kk}<0$), solute-vacancy drag ($L_{kV}/L^{V}_{kk}>0$) or vacancy wind effect, and solute-SIA binding can be incorporated. While Manning’s relations [56] for concentrated multicomponent alloy were used in this work to parameterize $L^{\nu}_{kl}$, one could utilize analytic methods (e.g., the self-consistent mean-field and the Green’s function) or atomistic calculations (e.g., kinetic Monte Carlo and molecular dynamics)  [13] to accurately parameterize the off-diagonal $L_{ij}$ coefficients for systems in which their contribution or solute-vacancy drag is deemed important [67].
@@
With the development of advanced, high-resolution compositional characterization, complex nanoscale segregation profiles across GBs, such as “W” and “M” contours, have been revealed, necessitating new models to explain experimental observations. The model formulation that we have developed enables a systematic investigation for TS without irradiation, RIS without TS, TS with radiation-enhanced diffusion (RETS), and RIS+RETS. Our results show that there are important characteristic differences in the segregation profiles depending on the active physical mechanism(s), temperature, and radiation damage rate. The model and its results enable a new means of interpreting experimental data. It also provides insights on how to structure experimental investigations into segregation mechanisms, given that the simulation results show qualitative differences in evolution with time and accumulated radiation damage. Thus, the interpretation of observed segregation behavior under irradiation is dependent upon the knowledge of the initial composition profile and the expected steady state, including the time and damage level to achieve it.

### v2

@@ 3 Parameterization
$P_{{\upsilon}}$ Point defect production rate $2\times 10^{-6}$ dpa/s
@@
[…]
$Q_{Fe}$ Activation energy for vacancy diffusion of Fe 2.89 eV [27]
$Q_{Cr}$ Activation energy for vacancy diffusion of Cr 2.88 eV [27]
$Q_{Ni}$ Activation energy for vacancy diffusion of Ni 2.86 eV [27]
$P_{Fe}$ Frequency factor for vacancy diffusion of Fe $3.8\times 10^{13}$ nm2/s [27]
$P_{Cr}$ Frequency factor for vacancy diffusion of Cr $5.5\times 10^{13}$ nm2/s [27]
$P_{Ni}$ Frequency factor for vacancy diffusion of Ni $1.5\times 10^{13}$ nm2/s [27]
@@ 3.1 Transport properties
Following Yang et al. [27], the Onsager coefficients are calculated from the partial diffusivities $d_{k{{\upsilon}}}$ by using Manning’s relations for concentrated multicomponent alloys [56]:
$L^{{\upsilon}}_{kl}=\left(\frac{c_{k}c_{{\upsilon}}d_{k{{\upsilon}}}}{RT/V_{m}}\right)\left(\delta_{kl}+\frac{2c_{l}d_{l{{\upsilon}}}}{M_{o}\sum_{j}c_{j}d_{j{{\upsilon}}}}\right),$ (18)
@@
Since the vacancy activation energies $Q_{k}$ are derived from experimental tracer diffusivity data, the vacancy migration energy for diffusion is given by $E^{m}_{Vk}=Q_{k}-E^{f}_{V}$. $Q_{k}$ is found to lie in the 2.86–2.89 eV range, and $E^{f}_{V}=1.8$ eV is chosen based on a first-principles study on dilute FCC Fe-Cr-Ni [58]. This yields an effective migration energy in the range of 1.05 to 1.1 eV. With this approach, $E^{f}_{V}$ is an uncertain parameter that must be more accurately estimated for a concentrated austenitic SS composition. The vacancy diffusivity at 500 C is $D_{V}\approx 5.5\times 10^{5}$ nm2/s and the partial diffusion coefficient ratios are $d_{CrV}/d_{FeV}=1.62$ and $d_{NiV}/d_{FeV}=0.67$. These ratios reflect the faster diffusivity of Cr and the slower diffusivity of Ni in FCC Fe. The SIA parameters in the model are quite uncertain, since they are not experimentally accessible and are challenging to determine via atomistic techniques. $E^{m}_{I}$ = 0.9 eV, as chosen by Yang et al. [27], resulted in a low diffusivity for SIA in comparison to the vacancy. Therefore, we chose a lower value of $E^{m}_{I}=0.6$ eV [27], yielding $D_{I}\approx 1.3\times 10^{7}$ nm2/s at 500 C—an order of magnitude higher than $D_{V}$. With regards to SIA binding, first-principles studies of SIA in dilute [58] and concentrated [59] FCC Fe–Ni–Cr indicate an attractive interaction between SIAs and Cr, a repulsive interaction with Ni, and an ideal interaction with Fe. Thus, the positive value for $E^{b}_{CrI}-E^{b}_{FeI}$ models favorable transport of Cr via SIAs, whereas the negative value for $E^{b}_{NiI}-E^{b}_{FeI}$ models unfavorable transport of Ni via SIAs. The resulting binding factors for the SIA partial diffusivities are ${\beta}_{CrI}=1.37$, ${\beta}_{NiI}=0.84$, and ${\beta}_{FeI}=0.94$. The ratios of SIA partial diffusion coefficients are $d_{CrI}/d_{FeI}=1.63$ and $d_{NiI}/d_{FeI}=0.89$. At lower temperatures ($<500^{\circ}$ C), the increased importance of Cr-SIA binding means that Cr transport via SIA is expected to become more important (i.e. $d_{CrI}/d_{FeI}>d_{CrV}/d_{FeV}$). All simulation of RIS performed in this paper employ the preferential solute-SIA binding discussed above. A comparison of these results with the case without solute-SIA binding is performed in Sec. S4 of the Supplementary Material.
@@ 3.3 Point defect reaction rates
The reaction rate coefficients in Eq. 12 and 14 are defined as follows. We set $P_{V}=P_{I}=2\times 10^{-6}$ dpa/s representing conditions of neutron or proton irradiation [23]. For the present work, we do not consider the formation and effects of clusters in the collision cascade and thus neglect the resulting point defect production bias. The recombination reaction coefficient is given by $R_{VI}=4\pi r_{\circ}(D_{V}+D_{I})/V_{a}$, with $r_{\circ}$ being the recombination radius. The point defect diffusivities are derived from the partial diffusivities in Sec. 3.1 as $D_{{\upsilon}}=(c_{Ni}d_{Ni{{\upsilon}}}+c_{Cr}d_{Cr{{\upsilon}}}+c_{Fe}d_{Fe{{\upsilon}}})/f_{\circ}$.
The sink strength of dislocation is given by $k^{2}_{{{\upsilon}},b}=\rho_{b}Z_{{\upsilon}}$, where $Z_{{\upsilon}}$ is the sink efficiency of dislocations for absorption of point defects. Realistically, $Z_{{\upsilon}}$ is a function of the dislocation density and character [63]. Relations for $Z_{{\upsilon}}$ based on calculations of the elastic interaction between dislocation configurations and the point defect have revealed a bias for SIA absorption (i.e., $Z_{I}>Z_{V}$), and this bias increases with dislocation density. For the sake of simplicity, we choose dislocation-density-independent values of $Z_{I}=1.2$ and $Z_{V}=1$, amounting to a bias factor $(Z_{I}-Z_{V})/Z_{V}$ of 20% [54]. We note that estimates of bias factors ranging from 2 to 30% have been reported using different modeling techniques and the discrepancy between these have been discussed in Ref. [64]. The choice of $Z_{I}=1.2$ therefore represents a strong bias.
@@ 3.4 Equilibrium PF properties
For the AM microstructure, per the literature [9, 10], a CW width of $\delta=100$ nm is considered. We choose a dislocation density of $\rho_{b,c}=10^{14}$ m-2 in the cell, and set $\rho_{b,w}$ in the CW to be 5–15 times higher. For $\rho_{b,w}=10^{15}$ m-2 distributed uniformly across the CW width, the dislocation energy of the CW can be estimated as $\frac{1}{2}Gb^{2}\rho_{c,w}\delta\approx 1$ J/m2, where $G=77$ GPa is the shear modulus. To realize a stationary CW of $\delta=100$ nm, we set the PF model parameters as $m_{0}=7.5\times 10^{6}$ J/m3 and $\kappa=7.5\times 10^{-8}$ J/m from Sec. S6 of the Supplementary Material. The interpolation scheme to determine the total atomic thermodynamic factors in Eq. 12 and 14 from the phase-specific atomic thermodynamic factors obtained above is provided in Sec. S3 of the Supplementary Material. The thermodynamic factors for the point defect species are simply given by ${\theta}_{{{\upsilon}}{{\upsilon}}}=(RT/V_{m})/c_{{\upsilon}}$.
@@ 4 Implementation
The model is implemented using the open-source MOOSE (Multiphysics Object Oriented Simulation Environment) framework. Weak forms of the partial differential equations are spatially discretized via the finite element method. For time integration, the implicit second-order backward differentiation method is used to obtain the system of nonlinear equations at each time step. These are then solved using Newton’s method. Mesh elements with linear Lagrange shape functions are used for the nonlinear and auxiliary variables. 2D domains are meshed with four-node quadrilateral elements. For both 1D and 2D simulations, a uniform mesh with adaptive time stepping (see Sec. S7 for details on the IterationAdaptiveDT scheme and the results of testing the parameters), a nonlinear relative tolerance of $10^{-8}$, and a nonlinear absolute tolerance of $10^{-10}$ are employed. The order parameters for the 2D polycrystal and 2D CW microstructures (Fig. 2) are generated using the Voronoi tessellation and grain growth algorithm implemented in MOOSE [66].
@@ 4.1 GB segregation
Simulations of GB segregation are performed in 1D and 2D with periodic boundary conditions. For the PF implementations, Eqs. 1212, and 13 are solved. For a GB width of $\delta=1$ nm, a mesh element size of $0.1$ nm is chosen, based on a mesh convergence test (i.e. separate simulations with different mesh element sizes were performed to identify the mesh element size at which the numerical solutions converge). The different PF simulation methods adopted in this work are as follows:
@@
In addition to the above PF simulations, we performed sharp interface simulations (see A and B) for RIS at GBs by employing Eq. 12 with boundary conditions. For this, we simply omit the PF sink term $-k^{2}_{g}D_{{\upsilon}}(c_{{\upsilon}}-c^{e}_{{\upsilon}})g_{\text{sink}}$ from Eq. 12 and impose a Dirichlet boundary condition (DBC) or Robin boundary condition (RBC) for the point defect concentration variables at the edges of the domain defining the GB. The DBC sets $c_{{\upsilon}}=c^{e}_{{\upsilon}}$ at the GB, thus describing an ideal sink. The RBC is applied as a flux normal to the GB given by $J_{{\upsilon}}.\hat{n}=D_{{\upsilon}}\alpha^{-1}(c_{{\upsilon}}-c^{e}_{{\upsilon}})$, where $\alpha$ is a characteristic length for point defect absorption  [65]. The effect of this parameter on the non-ideal GB sink behavior is studied in  B. The different sharp-interface simulations of RIS at GBs performed in this work are:
@@ 5 Results
We first present the results of GB segregation in 1D and 2D systems as a function of irradiation dose and temperature. We examine four different GB segregation mechanisms: RIS without TS, TS without RIS, radiation-enhanced TS (RETS) without RIS, and RIS in combination with RETS (RIS+RETS). The effects of grain size and dislocation bias on RIS are then presented. The results obtained through verification and comparison of the 1D PF RIS simulations against sharp-interface simulations are provided in A and B. Next, we present the results of RIS at dislocation CWs in 1D and 2D cells. The effects of cell size, CW density, initial microsegregation, and GB sink are examined. Results of point defect concentrations are provided in Sec. S8 of the Supplementary Material.
@@ 5.1.1 Segregation profiles
Concentration profiles for the RIS mechanism are plotted for different doses in Fig. 3(a). Ni enrichment (inverted “V”-shaped profile) and Cr depletion (“V”-shaped profile) are observed, with wide segregation widths of over 100 nm on either side of the GB center. Significant changes to RIS were found to occur between 0.01 dpa (1 hour) and 10 dpa (60 days). Concentration profiles for the RETS mechanism are plotted in Fig. 3(b). Unlike RIS, RETS is seen to evolve rapidly and reach equilibrium at just after 0.01 dpa (1 hour). Moreover, Ni depletion and Cr enrichment, which are opposite to their behaviors during RIS, are observed. In further contrast, RETS shows sharp profiles that only extend to about 0.5 nm on either side of the GB center. Next, Fig. 4 shows profiles from the combined RIS+RETS simulation. Here, two different cases are shown, with TS being strong ($\sigma=0.8$) in one case and moderate ($\sigma=0.9$) in the other. While the low-magnification plots of the profiles (see Sec. S8 in the Supplementary Material) resemble the monotonic RIS profiles, the high-magnification plots (Figs. 4(a) and 4(b)) clearly show non-monotonic variation close to the GB. For the case with strong TS, a slight Cr enrichment persists at the GB center, even at 10 dpa. The profiles resemble a “W” shape for Cr and an “M” shape for Ni. The non-monotonic profiles are less pronounced for the case of weak TS. Note that for the sake of simplicity, the local GB sink strength $k^{2}_{g}$ was kept the same and only $\sigma$ was varied; however, in a more realistic scenario, both $k^{2}_{g}$ and $\sigma$ would vary as a function of the GB structure.
@@ 5.1.2 Dose and temperature dependence
However, TS without irradiation shows lower segregation at the end of the simulated 60 days than that observed with RETS below 450 C, indicating that equilibrium TS is not achieved during that time frame due to the extremely slow diffusivity at lower temperatures. In contrast, RIS varies strongly with temperature and peaks at around 500 C. Negligible RIS is observed above 750 C due to thermal back diffusion. Conversely, below 500 C, increasing contributions from recombination and preferential SIA diffusivity (see Sec. S4 in the Supplementary Material) lead to decreasing RIS (here, SIAs are mobile but vacancies have reduced mobility; thus, recombination is promoted, resulting in a reduced loss of vacancies to sinks). Compared to RIS alone, RIS+RETS for $\sigma=0.8$ (strong TS) shows lower Ni enrichment at intermediate temperatures, while a transition to Ni depletion is observed below 400 C and above 650 C. For RIS+RETS of Cr, RETS dominates across the temperature range, resulting in an effective Cr enrichment at the GB center. However, at 0.5 nm from the GB center, there is negligible influence from RETS; therefore, the temperature trend follows that of RIS. Finally, we note that for $\sigma=0.95$ (weak TS, see Sec. S8 of the Supplementary Material), the RIS+RETS results are very close to that of RIS, even at the GB center.
@@ 6.1 PF model for GB segregation
A multi-order-parameter PF model was developed that describes both RIS and RETS in multicomponent polycrystals. The formulation also preserves the complete set of Onsager coefficients from the Onsager force-flux relations, thus allowing non-ideal kinetics ($L_{kV}/L^{{\upsilon}}_{kk}\neq-1$) to be captured. Therefore, RIS mechanisms of solute-vacancy exchange or inverse Kirkendall effect ($L_{kV}/L^{V}_{kk}<0$), solute-vacancy drag ($L_{kV}/L^{V}_{kk}>0$) or vacancy wind effect, and solute-SIA binding can be incorporated. While Manning’s relations [56] for concentrated multicomponent alloy were used in this work to parameterize $L^{{\upsilon}}_{kl}$, one could utilize analytic methods (e.g., the self-consistent mean-field and the Green’s function) or atomistic calculations (e.g., kinetic Monte Carlo and molecular dynamics)  [13] to accurately parameterize the off-diagonal $L_{ij}$ coefficients for systems in which their contribution or solute-vacancy drag is deemed important [67].
@@
With the development of advanced, high-resolution compositional characterization, complex nanoscale segregation profiles across GBs, such as “W” and “M” contours, have been revealed, necessitating new models to explain experimental observations. The model formulation that we have developed enables a systematic investigation for TS without irradiation, RIS without TS, TS with radiation-enhanced diffusion (RETS), and RIS+RETS. Our results show that there are important characteristic differences in the segregation profiles depending on the active physical mechanism(s), temperature, and radiation damage. The model and its results enable a new means of interpreting experimental data. It also provides insights on how to structure experimental investigations into segregation mechanisms, given that the simulation results show qualitative differences in evolution with time and accumulated radiation damage. Thus, the interpretation of observed segregation behavior under irradiation is dependent upon the knowledge of the initial composition profile and the expected steady state, including the time and damage level to achieve it.